# Navigating the Reality Gap: Privacy-Preserving On-Device Continual Adaptation of ASR for Clinical Telephony

### v3

@@ A.1 Hyperparameter Configurations
A.2 Algorithm and Buffer Management
The technical contribution of our framework lies in the efficient integration of LoRA, linearized EWC, and prioritized experience replay. Algorithm A.2 details the localized adaptation loop, and Algorithm A.2 describes our multi-domain buffer management strategy.
Algorithm 1: LoRA-based Hybrid Adaptation Loop Require: Base model $\theta_{base}$, Replay buffer $\mathcal{B}$, Regularization $\lambda$ 1: Initialize LoRA adapters $\theta_{0}\subset\theta_{base}$ and importance $F\leftarrow\mathbf{0}$ 2: for segment $k=1,\dots,K$ do
3:
 Receive clinical stream $\mathcal{D}_{k}$
[…]
5:Adaptation Step:
  $\theta_{k}\leftarrow\text{argmin}_{\theta}\mathcal{L}_{CTC}(\mathcal{D}_{train})+\frac{\lambda}{2}\sum F_{i}(\theta_{k}-\theta_{k-1}^{*})^{2}$
6:Importance Estimation (Diagonal Approx.):
  $F_{new}\leftarrow\frac{1}{N}\sum_{x\in\mathcal{D}_{train}}|\nabla_{\theta}\mathcal{L}_{CTC}(x)|$ Absolute Fisher
7:Asynchronous Consolidation:
  $F\leftarrow\frac{F\times(k-1)+F_{new}}{k}$ Segment-wise Average
[…]
10: end for
Algorithm 2: Multi-Domain Buffer Management (Prioritization) Require: Segment data $\mathcal{D}_{k}$, Buffer $\mathcal{B}$, Loss threshold $\tau$ 1: Compute $\mathcal{L}_{inst}(x)$ for all instances $x\in\mathcal{D}_{k}$ 2: Hard Example Mining (Prioritization):
3: Buffer Update:
$\mathcal{B}_{gv}\leftarrow\text{Sample}(60\%\text{ from }\mathcal{S}_{hard}\cup 40\%\text{ Random})$
$\mathcal{B}_{gen}\leftarrow\text{Sample}(300\text{ Balanced samples from }\text{Kathbath})$ 4: $\mathcal{B}\leftarrow\mathcal{B}_{gv}\cup\mathcal{B}_{gen}$
A.3 Computational Efficiency and Hardware Setup
A core goal of our work is to prove that high-performance, privacy-preserving adaptation is possible on standard mobile workstations rather than centralized server farms. We evaluated our pipeline on two laptop configurations:
  1. 1.
    Intel i7-13700H (14 cores) CPU + NVIDIA RTX 4050 (35W) GPU. Training time averaged 25–30 minutes per data segment.
  2. 2.
    Intel i5-12500H (12 cores) CPU + NVIDIA RTX 3050 (70W) GPU. Training time was more variable but stayed within 50–60 minutes per segment.
These benchmarks demonstrate the technical feasibility of localized, self-improving ASR systems in environments where high-end compute clusters are unavailable.
Appendix B Detailed Experimental Results
B.1 Adaptation Dynamics: Naive and Single-Domain Replay
Figure 4: Detailed adaptation dynamics for Naive (V1, V1.1) paradigms.
Figure 5: Detailed adaptation dynamics for Single-Domain Replay (V2, V2.1) paradigms.
B.2 Adaptation Dynamics: Multi-Domain Replay
Figure 6: Detailed adaptation dynamics for Multi-Domain Replay (V3, V3.1) paradigms.
B.3 Adaptation Dynamics: EWC and Hybrid Strategies
Figure 7: Detailed adaptation dynamics for EWC strategies with high/medium regularization (V4.3, V4.4).
Figure 8: Detailed adaptation dynamics for optimal EWC (V4.5) and Hybrid ER+EWC strategies (V5, V5.1).
B.4 Catastrophic Forgetting Analysis
The preservation of foundational general-domain knowledge (Kathbath) is evaluated across all experimental paradigms. Figures 9, 10, 11, and 12 illustrate the performance stability on the source domain during continuous adaptation.
Figure 9: Catastrophic forgetting analysis for Conservative strategies (V1, V2, V3).
Figure 10: Catastrophic forgetting analysis for Aggressive Warmup strategies (V1.1, V2.1, V3.1).
Figure 11: Impact of EWC regularization ($\lambda$) on Catastrophic Forgetting (V4.3, V4.4, V4.5).
Figure 12: Catastrophic forgetting analysis for Hybrid ER+EWC strategies (V5, V5.1).
B.5 Numerical Stability and Gradient Analysis
A critical challenge encountered during the development of the EWC-based paradigms (V4 series) was the phenomenon of gradient explosion. In early iterations of the V4.1 paradigm, gradients used for Fisher Information consolidation were not properly detached from the computation graph. This caused the regularization term to recursively accumulate gradient histories across segments, leading to a computational overhead and gradient norms exceeding 1000. Even after fixing the detaching logic, the standard quadratic Fisher importance in V4.1 remained highly sensitive to noisy data, frequently dominating the optimization objective. This numerical instability caused the model parameters to “freeze,” preventing the system from learning the target domain features.
@@ B.5 Numerical Stability and Gradient Analysis
Figure 13: Gradient stability analysis comparing Standard EWC (V4.1), Naive Baseline (V1.1), and our proposed Hybrid L-EWC (V5.1). The log-scale plot demonstrates how the linearized importance estimation in L-EWC prevents the gradient explosion seen in standard quadratic formulations.
As shown in Figure 13, our proposed Linearized EWC (L-EWC) strategy successfully stabilizes the gradient norm throughout the 24 adaptation segments. By utilizing absolute gradient values for importance estimation, we ensure that the optimization trajectory remains healthy, with gradient magnitudes comparable to the naive baseline while still enforcing the necessary stability constraints to prevent forgetting.
Appendix C Dataset Characteristics
The Gram Vaani dataset serves as a rigorous proxy for rural clinical environments due to its telephonic acquisition (originally 8kHz upsampled to 16kHz) and focus on medical/agricultural discussions. Table 4 provides a breakdown of the dataset characteristics.
Metric Value
Total Duration 103.2 Hours
Number of Segments ($k$) 24
Samples per Segment $\approx$ 1,600
Avg. Duration per Sample 9.4 Seconds
Acoustic Condition Noisy Telephonic (8kHz up)
Primary Dialect Rural Hindi (various)

### v4

@@ Appendix C Dataset Characteristics
Configuration CPU GPU Training Time (per segment)
Config A (High-End) Intel i7-13700H (14 cores) NVIDIA RTX 4050 (35W) 25–30 minutes
Config B (Mid-Range) Intel i5-12500H (12 cores) NVIDIA RTX 3050 (70W) 50–60 minutes
Table 5: Hardware configurations and training time benchmarks for on-device adaptation.
Figure 4: Naive Baseline Dynamics (V1, V1.1): Showing rapid initial adaptation but significant volatility, highlighting the need for regularization.
Figure 5: Single-Domain Replay Dynamics (V2, V2.1): Incorporating target-domain replay buffers stabilizes the learning trajectory compared to naive fine-tuning.
Figure 6: Multi-Domain Replay Dynamics (V3, V3.1): Balancing target and general domain samples in the replay buffer offers a tradeoff between plasticity and stability.
Figure 7: EWC Regularization (V4.3, V4.4): High regularization ($\lambda$) dampens plasticity, while moderate values allow for adaptation, though standard EWC shows instability.
Figure 8: Optimal Hybrid Strategies (V5, V5.1): Combining Linearized EWC with Replay (V5.1) achieves the lowest final WER with stable convergence.
Figure 9: Forgetting Analysis (Conservative): Baseline strategies (V1-V3) show varying degrees of knowledge retention on the Kathbath dataset.
Figure 10: Forgetting Analysis (Aggressive): Increased learning rates in V1.1-V3.1 accelerate adaptation but risk higher catastrophic forgetting.
Figure 11: EWC Forgetting Analysis: Stronger regularization ($\lambda$) effectively reduces forgetting but may hinder adaptation speed.
Figure 12: Hybrid Forgetting Analysis (V5, V5.1): The hybrid approach demonstrates the best balance, minimizing forgetting while maintaining plasticity.
@@
Figure 13: Gradient Stability Analysis: Comparing Standard EWC (V4.1), Naive Baseline (V1.1), and our proposed Hybrid L-EWC (V5.1). The log-scale plot demonstrates how the linearized importance estimation in L-EWC prevents the gradient explosion seen in standard quadratic formulations.